OpenStudy (anonymous):

given sin sigma = -4/5 and tan sigma>0 , what are the remaining functions of trig values?

OpenStudy (anonymous):

MUSTARD

OpenStudy (anonymous):

The negative sign indicates the angle will be in quadrant 3 or 4. The positive tan indicates the angle must be in 3; since cos must also be negative. cos =\[ \sqrt{1-\sin ^{2}\sigma}\] = -3/5 The rest you can calculate from there.